# assumeAssumption

### v1

@@ 2.5 O-minimal structures
where $f_{1},\ldots,f_{q}$ are polynomials with real coefficients. We adopt [41, Definition p. 503-506] below.
Definition 1.
@@
A key property of univariate definable functions is that they satisfy the monotonicity theorem [41, 4.1]. It states that on bounded open intervals, for any $p\in\mathbb{N}$ there exist finitely many open subintervals where the function is $C^{p}$ and either constant or strictly monotone. The extension of the monotonicity theorem to multivariate functions is the cell decomposition theorem [41, 4.2], which is used to prove the definable Morse-Sard theorem [8, Corollary 9]. It asserts that that lower semi-continuous definable functions have finitely many Clarke critical values.
@@ 3.1 Definition of flatness
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is constant near $x\in\mathbb{R}^{n}$ if it is constant on a neighborhood of $x$. Otherwise, $f$ is nonconstant near $x$, which happens iff $\accentset{\circ}{f}(x,r)>0$ for all $r>0$ iff $x\notin\mathrm{int}[f=f(x)]$. Given $x$, $f$ is constant iff $\overline{f}(x,\ell)=\infty$ for all $\ell>0$, and $f$ is continuous at $x$ iff $\overline{f}(x,\ell)>0$ for all $\ell>0$. A set $X\subset\mathbb{R}^{n}$ is nowhere dense if its closure has empty interior. If $f$ is continuous and $\alpha\in\mathbb{R}$, then $f$ is nonconstant near every point in $f=\alpha]$ iff $f=\alpha]$ is nowhere dense.
A function $\varphi:S\to\overline{\mathbb{R}}$ where $S\subset\mathbb{R}$ is increasing (resp. strictly increasing) if $\varphi(s)\leq\varphi(t)$ (resp. $\varphi(s)<\varphi(t)$) for all $s,t\in S$ such that $s<t$. It is monotone (resp. strictly monotone) if either $\varphi$ or $\varphi$ is increasing (resp. strictly increasing). Given $x\in\mathbb{R}^{n}$, let $\accentset{\circ}{f}_{x}=\accentset{\circ}{f}(x,\cdot)$ and $\overline{f}_{x}=\overline{f}(x,\cdot)$, both of which are increasing.
Fact 2.
@@
We now gather several useful facts when $f$ is continuous (i.e., Facts 3-8).
Fact 3.
@@
˜4 and ˜5 actually hold without continuity if one uses strict relations, including in the definition of $\accentset{\circ}{f}$ and $\overline{f}$, with the convention that $\accentset{\circ}{f}(x,0)=\overline{f}(x,0)=0$.
Fact 6.
@@
(i) $\Longrightarrow$ (ii) Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})>0$ and $\ell\in(0,\overline{\ell}]$. By ˜4,
@@
(ii) $\Longrightarrow$ (i) Since $f$ is nonconstant, $\infty>\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})>0$ after possibly reducing $\overline{\ell}$. Let $\overline{r}=\accentset{\circ}{f}(y,\overline{\ell})/2>0$ and $r\in(0,\overline{r}]$. Since $\overline{f}(x,\ell)\geq\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})=2\overline{r}>r$ for all $\ell\in(\overline{\ell},\infty)$, by ˜5,
@@
Since $\accentset{\circ}{f}$ is continuous, one has equality in ˜4, namely, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)=\ell\}$. This holds by the intermediate value theorem as $\accentset{\circ}{f}(x,0)=0$. In contrast, since $\overline{f}$ is merely lower semicontinuous, equality does not necessarily hold in ˜5 (think of the Cantor function). Likewise, $\accentset{\circ}{f}_{x}$ need not be strictly increasing. To obtain such properties and others, more assumptions are needed.
Assumption 1.
@@
(i) & (ii) Continuity of $\accentset{\circ}{f}_{x}$ follows from ˜7. Clearly $\accentset{\circ}{f}_{x}$ is increasing on $0,\overline{r}]$. It is in fact strictly increasing. Indeed, suppose $\accentset{\circ}{f}_{x}$ is constant near $r\in(0,\overline{r})$. If $\accentset{\circ}{f}(x,r)=0$, then $f$ is constant near $x$, violating our assumption. Otherwise, by continuity there exists $y\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(y)-f(x)|$, which must be a local maximum of |ff(x)||f-f(x)|. Since $\accentset{\circ}{f}(x,r)>0$, $y$ is either a local minimum or a local maximum of $f$, and hence $0\in\partial f(y)=\mathrm{co}\hskip 0.85358pt\partial f(y)=\overline{\partial}f(y)$ or $0\in\partial(-f)(y)\subseteq\overline{\partial}(-f)(y)=-\overline{\partial}f(y)$ by Fermat’s rule [27, Theorem 10.1]. Since $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}}(x)$, we have $f(y)=f(x)$. It follows that $\accentset{\circ}{f}(x,r)=0$, a contradiction.
By ˜4 and the discussion below ˜8, $\accentset{\circ}{f}_{x}(\overline{f}_{x}(\ell))=\ell$ for all $\ell\in[0,\overline{\ell}]$. Thus for all $r\in[0,\overline{r}]$, $\accentset{\circ}{f}_{x}((\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r))=(\accentset{\circ}{f}_{x}\circ\overline{f}_{x})(\accentset{\circ}{f}_{x}(r))=\accentset{\circ}{f}_{x}(r)$ and $(\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r)=r$. It follows that $\overline{f}_{x}$ is continuous and strictly increasing.
@@
The first point follows from ˜6, while the second follows from Facts 2-7, Proposition˜1, and the definable Morse-Sard theorem [8, Corollary 9]. ∎
@@ 3.2 Curve selection
    There exist $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that for all $r\in[0,\overline{r}]$, $\gamma(r)=x+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ and $v(r)\in\overline{\partial}f(\gamma(r))$.
@@
    If $x$ is a local minimum of $f$, then $\lambda(r)>0$ and $v(r)\in-\partial(-f)(\gamma(r))$ for all $r\in(0,\overline{r}]$ after possibly reducing $\overline{r}$. If in addition, $f$ is regular near $x$, then $v(r)\in\partial f(\gamma(r))$ for all $r\in(0,\overline{r}]$.
@@
(i) Let $(r,\ell)\in\mathrm{gph}\accentset{\circ}{f}_{x}$. There exists $\gamma(r)\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(\gamma(r))-f(x)|$. By [36, Theorem 8.15], $0\in\partial|f-f(x)|(\gamma(r))+N_{\overline{B}_{r}(x)}(\gamma(r))$. If $r=0$, then let $\lambda(0)=0$ and $v(0)=0$. Otherwise, there exist $w(r)\in\partial(s(r)f)(\gamma(r))\subseteq\mathrm{co}\hskip 0.85358pt\partial(s(r)f)(\gamma(r))=\overline{\partial}(s(r)f)(\gamma(r))=s(r)\overline{\partial}f(\gamma(r))$ with $s(r)=\mathrm{sign}(f(x)-f(\gamma(r)))$ and $\mu(r)\geq 0$ such that $w(r)+\mu(r)(\gamma(r)-x)=0$. If $\mu(r)=0$, then $w(r)=0$ and so $f(\gamma(r))$ is a Clarke critical value of $f$. By assumption $f(\gamma(r))=f(x)$ and thus $\accentset{\circ}{f}_{x}(r)=0$, i.e., $r=0$, a contradiction. Thus $\mu(r)>0$. We may hence define $\lambda(r)=-s(r)/\mu(r)$ and $v(r)=s(r)w(r)$.
(ii) When $x$ is a local minimum of $f$, s(r)=1s(r)=-1. By [36, Theorem 9.16],
$\lim_{\tau\searrow 0}\frac{f(\gamma(r)+\tau w)-f(\gamma(r))}{\tau}=\lim_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\gamma(r)+\tau w^{\prime})-f(\gamma(r))}{\tau}=\max_{v\in\partial f(\gamma(r))}\langle w,v\rangle.$
Since the directional derivative is nonpositive for any direction $w$ such that $\langle w,\gamma(r)-x\rangle<0$, this also holds for any direction $w$ such that $\langle w,\gamma(r)-x\rangle=0$. Given $v\in\partial f(\gamma(r))$, let $w=v|\gamma(r)-x|^{2}-\langle v,\gamma(r)-x\rangle(\gamma(r)-x)$. Since $\langle w,\gamma(r)-x\rangle=0$, it follows that $\langle w,v\rangle=|v|^{2}|\gamma(r)-x|^{2}-\langle v,\gamma(r)-x\rangle^{2}\leq 0$. There is equality in the Cauchy-Schwarz inequality, so there exists $\mu\in\mathbb{R}$ such that $v=\mu(\gamma(r)-x)$. Since $\langle x-\gamma(r),\gamma(r)-x\rangle<0$, we have $\langle x-\gamma(r),v\rangle=-\mu|\gamma(r)-x|^{2}\leq 0$ and so $\mu\geq 0$. Since $\partial f(\gamma(r))\neq\emptyset$ by [36, Theorem 9.13] and $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}(x)}$, there exists $v(r)\in\partial f(\gamma(r))\setminus\{0\}$ and it satisfies $\gamma(r)-x=\lambda(r)v(r)$ for some $\lambda(r)>0$.
@@
(iii) When $f$ is definable, $0,\overline{r}]\ni r\rightrightarrows\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ is definable, there exists a definable selection $\gamma$ [41, 4.5]. Consequently $0,\overline{r}]\ni r\rightrightarrows\{v\in\overline{\partial}f(\gamma(t)):\exists\lambda\in\mathbb{R}:\gamma(t)=x+\lambda v\}$ is definable, yielding a definable selection $v$. Finally, $\lambda(r)=\langle\gamma(r)-x,v(r)\rangle/|v(r)|^{2}$ is definable. The monotonicity theorem ensures $\gamma,\lambda,v$ are $C^{k}$ for any $k\in\mathbb{N}$ on $(0,\overline{r}]$ after possibly reducing $\overline{r}$. Also $|\gamma_{i}(r)|\leq|\gamma(r)|=r$ so $|\gamma^{\prime}_{i}(r)|\leq 1$ on $(0,\overline{r}]$ after possibly reducing $\overline{r}$ and $|\gamma^{\prime}(r)|=\sqrt{|\gamma^{\prime}_{1}(r)|^{2}+\cdots+|\gamma^{\prime}_{n}(r)|^{2}}\leq\sqrt{n}$. Hence $\lim_{r\searrow 0}\gamma^{\prime}(r)$ exists. By the mean value theorem, $\gamma_{i}^{\prime}(0)=\lim_{r\searrow 0}(\gamma_{i}(r)-x_{i})/r=\lim_{r\searrow 0}\gamma_{i}^{\prime}(r)$. Finally, one can take $v(0)=\lim_{r\searrow 0}v(r)\in\overline{\partial}f(\gamma(0))$, where the limit exists because $v(r)$ is bounded by [36, Theorem 9.13] and [4, Proposition 3 p. 42].
@@ 3.3 Calculus rules
Since $f$ is Lipschitz continuous near $x$, by [36, Theorem 9.13] $\partial f(x)$ is nonempty and compact, and $\mathrm{lip}f(x)=\max\{|v|:v\in\partial f(x)\}$. Since $f$ is Lipschitz continuous and regular near $x$, by [36, Theorem 9.16] we have $\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau=\max\{\langle u,v\rangle:v\in\partial f(x)\}$. Thus
@@
for $r>0$ near 0. On the other hand, let $\overline{u}\in\arg\max_{|u|=1}\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau$. By Lemma˜2, we have $\accentset{\circ}{f}(x,r)\geq|f(x+r\overline{u})-f(x)|=\mathrm{lip}f(x)r+o(r)$ for $r\geq 0$ near 0. ∎
Without regularity, Proposition˜3 may fail.
Example 1.
@@
Proposition˜3 and Proposition˜4 admit the following corollaries.
Corollary 1.
@@
and in particular $\accentset{\circ}{f}(\overline{x},r)/r\leq\accentset{\circ}{f}(x,r)/r$ for all $r\in(0,\overline{r}]$. Passing to the limit as $r\searrow 0$ yields $\mathrm{lip}f(\overline{x})\leq\mathrm{lip}f(x)$ by Proposition˜3.
(ii) There exists a neighborhood $U$ of $\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(\overline{x})<\mathrm{lip}f(x)$ for all $x\in U\setminus\{\overline{x}\}$. By Proposition˜3, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in U$ after possibly reducing $U$. Thus, for all $x\in U\setminus\{\overline{x}\}$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{x},r)<\accentset{\circ}{f}(x,r)$ for all $r\in(0,\overline{r}]$. ∎
Corollary 2.
@@
One argues as for Corollary˜2 using Proposition˜4. ∎
@@
˜9 implies that $x\preceq y$ and $y\preceq x$ for all $x,y\in G\overline{x}$. By Proposition˜3, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in\mathbb{R}^{n}$. There exists a neighborhood $U$ of $G\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(x)<\mathrm{lip}f(y)$ for all $x\in G\overline{x}$ and $y\in U\setminus G\overline{x}$. Hence, for any such $x$ and $y$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(x,r)<\accentset{\circ}{f}(y,r)$ for all $r\in(0,\overline{r}]$. ∎
@@
Fact 10.
[36, Theorem 10.6] Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is locally Lipschitz regular and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. Then $f$ is regular and
@@
Remark 1.
Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is $C^{1}$ such that $\nabla g(0)=0$ and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is locally Lipschitz. The Jacobian chain rule [11, Theorem 2.6.6] yields $\overline{\partial}f(x)=\overline{\partial}F(x)^{T}\nabla g(F(x))$. Hence, if $F(x)=0$, then $\overline{\partial}f(x)=\{0\}$. Since the Clarke subdifferential is a singleton, by [36, Theorem 9.18] $f$ is strictly differentiable [36, Definition 9.17] at $x$, $\nabla f(x)=0$, and $\mathrm{lip}f(x)=0$.
@@
Fact 12.
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be lower semicontinuous and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$. For all $\overline{x}\in\mathrm{dom}f$ and $g\in G$, $\partial f(g^{-1}\overline{x})=g^{T}\partial f(\overline{x})$.
@@
Applying the chain rule [36, Exercise 10.7] to $f(x)=f(gx)$ at $g^{-1}\overline{x}$ for any $g\in G\subseteq\mathrm{GL}(n,\mathbb{R})$ yields the result. ∎
Fact 13.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$. For all $\overline{x},v_{1},\ldots,v_{k}\in\mathbb{R}^{n}$, $f^{(k)}(g^{-1}\overline{x})(v_{1},\ldots,v_{k})=f^{(k)}(\overline{x})(gv_{1},\ldots,gv_{k})$.

### v2

@@ 2.5 O-minimal structures
where $f_{1},\ldots,f_{q}$ are polynomials with real coefficients. We adopt [49, Definition p. 503-506] below.
Definition 2.2.
@@
A key property of univariate definable functions is that they satisfy the monotonicity theorem [49, 4.1]. It states that on bounded open intervals, for any $p\in\mathbb{N}$ there exist finitely many open subintervals where the function is $C^{p}$ and either constant or strictly monotone. The extension of the monotonicity theorem to multivariate functions is the cell decomposition theorem [49, 4.2], which is used to prove the definable Morse-Sard theorem [9, Corollary 9]. It asserts that that lower semi-continuous definable functions have finitely many Clarke critical values.
@@ 3.1 Definition of flatness
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is constant near $x\in\mathbb{R}^{n}$ if it is constant on a neighborhood of $x$. Otherwise, $f$ is nonconstant near $x$, which happens iff $\accentset{\circ}{f}(x,r)>0$ for all $r>0$ iff $x\notin\mathrm{int}[f=f(x)]$. Given $x$, $f$ is constant iff $\overline{f}(x,\ell)=\infty$ for all $\ell>0$, and $f$ is continuous at $x$ iff $\overline{f}(x,\ell)>0$ for all $\ell>0$. A set $X\subseteq\mathbb{R}^{n}$ is nowhere dense if its closure has empty interior. If $f$ is continuous and $\alpha\in\mathbb{R}$, then $f$ is nonconstant near every point in $f=\alpha]$ iff $f=\alpha]$ is nowhere dense.
A function $\varphi:S\to\overline{\mathbb{R}}$ where $S\subseteq\mathbb{R}$ is increasing (resp. strictly increasing) if $\varphi(s)\leq\varphi(t)$ (resp. $\varphi(s)<\varphi(t)$) for all $s,t\in S$ such that $s<t$. It is monotone (resp. strictly monotone) if either $\varphi$ or $\varphi$ is increasing (resp. strictly increasing). Given $x\in\mathbb{R}^{n}$, let $\accentset{\circ}{f}_{x}=\accentset{\circ}{f}(x,\cdot)$ and $\overline{f}_{x}=\overline{f}(x,\cdot)$, both of which are increasing.
Fact 3.3.
@@
We now gather several useful facts when $f$ is continuous (i.e., Facts 3.4-3.9).
Fact 3.4.
@@
Fact˜3.5 and Fact˜3.6 actually hold without continuity if one uses strict relations, including in the definition of $\accentset{\circ}{f}$ and $\overline{f}$, with the convention that $\accentset{\circ}{f}(x,0)=\overline{f}(x,0)=0$.
Fact 3.7.
@@
(i) $\Longrightarrow$ (ii) Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})>0$ and $\ell\in(0,\overline{\ell}]$. By Fact˜3.5,
@@
(ii) $\Longrightarrow$ (i) Since $f$ is nonconstant, $\infty>\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})>0$ after possibly reducing $\overline{\ell}$. Let $\overline{r}=\accentset{\circ}{f}(y,\overline{\ell})/2>0$ and $r\in(0,\overline{r}]$. Since $\overline{f}(x,\ell)\geq\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})=2\overline{r}>r$ for all $\ell\in(\overline{\ell},\infty)$, by Fact˜3.6,
@@
Since $\accentset{\circ}{f}$ is continuous, one has equality in Fact˜3.5, namely, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)=\ell\}$. This holds by the intermediate value theorem as $\accentset{\circ}{f}(x,0)=0$. In contrast, since $\overline{f}$ is merely lower semicontinuous, equality does not necessarily hold in Fact˜3.6 (think of the Cantor function). Likewise, $\accentset{\circ}{f}_{x}$ need not be strictly increasing. To obtain such properties and others, more assumptions are needed.
Assumption 3.10.
@@
(i) & (ii) Continuity of $\accentset{\circ}{f}_{x}$ follows from Fact˜3.8. Clearly $\accentset{\circ}{f}_{x}$ is increasing on $0,\overline{r}]$. It is in fact strictly increasing. Indeed, suppose $\accentset{\circ}{f}_{x}$ is constant near $r\in(0,\overline{r})$. If $\accentset{\circ}{f}(x,r)=0$, then $f$ is constant near $x$, violating our assumption. Otherwise, by continuity there exists $y\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(y)-f(x)|$, which must be a local maximum of |ff(x)||f-f(x)|. Since $\accentset{\circ}{f}(x,r)>0$, $y$ is either a local minimum or a local maximum of $f$, and hence $0\in\partial f(y)=\mathrm{co}\hskip 0.85358pt\partial f(y)=\overline{\partial}f(y)$ or $0\in\partial(-f)(y)\subseteq\overline{\partial}(-f)(y)=-\overline{\partial}f(y)$ by Fermat’s rule [32, Theorem 10.1]. Since $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}}(x)$, we have $f(y)=f(x)$. It follows that $\accentset{\circ}{f}(x,r)=0$, a contradiction.
By Fact˜3.5 and the discussion below Fact˜3.9, $\accentset{\circ}{f}_{x}(\overline{f}_{x}(\ell))=\ell$ for all $\ell\in[0,\overline{\ell}]$. Thus for all $r\in[0,\overline{r}]$, $\accentset{\circ}{f}_{x}((\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r))=(\accentset{\circ}{f}_{x}\circ\overline{f}_{x})(\accentset{\circ}{f}_{x}(r))=\accentset{\circ}{f}_{x}(r)$ and $(\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r)=r$. It follows that $\overline{f}_{x}$ is continuous and strictly increasing.
@@
The first point follows from Fact˜3.7, while the second follows from Facts 3.3-3.8, Proposition˜3.11, and the definable Morse-Sard theorem [9, Corollary 9]. ∎
@@ 3.2 Curve selection
    There exist $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that for all $r\in[0,\overline{r}]$, $\gamma(r)=x+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$, $|\gamma(r)-x|=r$, and $v(r)\in\overline{\partial}f(\gamma(r))$.
@@
    If $x$ is a local minimum of $f$, then $\lambda(r)>0$ and $v(r)\in-\partial(-f)(\gamma(r))$ for all $r\in(0,\overline{r}]$. If in addition, $f$ is regular, then one may choose $v(r)\in\overline{\nabla}f(\gamma(r))$ for all $r\in(0,\overline{r}]$.
@@
(i) By Proposition˜3.11 (iii), there exists $\gamma(r)\in S_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(\gamma(r))-f(x)|$. The first-order optimality condition [44, Theorem 8.15] reads $0\in\partial(-|f-f(x)|)(\gamma(r))+N_{\overline{B}_{r}(x)}(\gamma(r))$. If $r=0$, then let $\lambda(0)=0$ and $v(0)\in\overline{\partial}f(x)\neq\emptyset$ by [44, Theorem 9.13]. Otherwise, there exist $w(r)\in\partial(s(r)f)(\gamma(r))\subseteq\mathrm{co}\hskip 0.85358pt\partial(s(r)f)(\gamma(r))=\overline{\partial}(s(r)f)(\gamma(r))=s(r)\overline{\partial}f(\gamma(r))$ with $s(r)=\mathrm{sign}(f(x)-f(\gamma(r)))$ and $\mu(r)\geq 0$ such that $w(r)+\mu(r)(\gamma(r)-x)=0$. If $\mu(r)=0$, then $w(r)=0$ and so $f(\gamma(r))$ is a Clarke critical value of $f$. By assumption $f(\gamma(r))=f(x)$ and thus $\accentset{\circ}{f}_{x}(r)=0$, i.e., $r=0$, a contradiction. Thus $\mu(r)>0$. We may hence define $\lambda(r)=-s(r)/\mu(r)$ and $v(r)=s(r)w(r)$.
(ii) When $x$ is a local minimum of $f$, s(r)=1s(r)=-1. Since $\gamma(r)$ is a local minimum of $g=-f+\delta_{B_{r}(x)}$, $dg(\gamma(r))\geq 0$ by Fermat’s rule [44, Theorem 10.1]. As $f$ is locally Lipschitz regular, by [44, Theorem 9.16] $f$ is semidifferentiable and $df(\cdot)=\sigma_{\partial f(\cdot)}$. Thus f-f is semidifferentiable and d(f)=dfd(-f)=-df. Since f-f is semidifferentiable, $dg=d(-f+\delta_{B_{r}(x)})=d(-f)+d\delta_{B_{r}(x)}$. Thus $dg(\gamma(r))=-\sigma_{\partial f(\gamma(r))}+\sigma_{N_{B_{r}(x)}(\gamma(r))}\geq 0$ and $\sigma_{\partial f(\gamma(r))}\leq\sigma_{N_{B_{r}(x)}(\gamma(r))}$, namely $\partial f(\gamma(r))\subseteq N_{B_{r}(x)}(\gamma(r))=\mathbb{R}_{+}(\gamma(r)-x)$. But $\partial f$ is nonempty convex compact valued by [44, Theorem 9.61], and so $\partial f(\gamma(r))=[a(r),b(r)](\gamma(r)-x)$ for some $0<a(r)\leq b(r)$ (recall that $0\notin\overline{\partial}f(\gamma(r))$). Also $\mathrm{co}\hskip 0.85358pt\overline{\nabla}f(\gamma(r))=\partial f(\gamma(r))$, hence $b(r)(\gamma(r)-x)\in\nabla f(\gamma(r))$. Now, simply let $v(r)=b(r)(\gamma(r)-x)$ and $\lambda(r)=1/b(r)>0$.
@@
(iii) When $f$ is definable, $0,\overline{r}]\ni r\rightrightarrows\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ is definable, so there exists a definable selection $\gamma$ [49, 4.5]. Consequently $0,\overline{r}]\ni r\rightrightarrows\{v\in\overline{\partial}f(\gamma(t)):\exists\lambda\in\mathbb{R}:\gamma(t)=x+\lambda v\}$ is definable, yielding a definable selection $v$. Finally, $\lambda(r)=\langle\gamma(r)-x,v(r)\rangle/|v(r)|^{2}$ is definable. The monotonicity theorem ensures $\gamma,\lambda,v$ are $C^{k}$ for any $k\in\mathbb{N}$ on $(0,\overline{r}]$ after possibly reducing $\overline{r}$. Also $|\gamma_{i}(r)|\leq|\gamma(r)|=r$ so $|\gamma^{\prime}_{i}(r)|\leq 1$ on $(0,\overline{r}]$ after possibly reducing $\overline{r}$ and $|\gamma^{\prime}(r)|=\sqrt{|\gamma^{\prime}_{1}(r)|^{2}+\cdots+|\gamma^{\prime}_{n}(r)|^{2}}\leq\sqrt{n}$. Hence $\lim_{r\searrow 0}\gamma^{\prime}(r)$ exists. By the mean value theorem, $\gamma_{i}^{\prime}(0)=\lim_{r\searrow 0}(\gamma_{i}(r)-x_{i})/r=\lim_{r\searrow 0}\gamma_{i}^{\prime}(r)$. Finally, one can take $v(0)=\lim_{r\searrow 0}v(r)\in\overline{\partial}f(\gamma(0))$, where the limit exists because $v(r)$ is bounded by [44, Theorem 9.13] and [4, Proposition 3 p. 42].
@@ 3.3 Calculus rules
Since $f$ is Lipschitz continuous near $x$, by [44, Theorem 9.13] $\partial f(x)$ is nonempty and compact, and $\mathrm{lip}f(x)=\max\{|v|:v\in\partial f(x)\}$. Since $f$ is Lipschitz continuous and regular near $x$, by [44, Theorem 9.16] we have $\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau=\max\{\langle u,v\rangle:v\in\partial f(x)\}$. Thus
@@
for $r>0$ near 0. On the other hand, let $\overline{u}\in\arg\max_{|u|=1}\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau$. By Lemma˜3.14, we have $\accentset{\circ}{f}(x,r)\geq|f(x+r\overline{u})-f(x)|=\mathrm{lip}f(x)r+o(r)$ for $r\geq 0$ near 0. ∎
Without regularity, Proposition˜3.15 may fail.
Example 3.16.
@@
Proposition˜3.15 and Proposition˜3.17 admit the following corollaries.
Corollary 3.18.
@@
and in particular $\accentset{\circ}{f}(\overline{x},r)/r\leq\accentset{\circ}{f}(x,r)/r$ for all $r\in(0,\overline{r}]$. Passing to the limit as $r\searrow 0$ yields $\mathrm{lip}f(\overline{x})\leq\mathrm{lip}f(x)$ by Proposition˜3.15.
(ii) There exists a neighborhood $U$ of $\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(\overline{x})<\mathrm{lip}f(x)$ for all $x\in U\setminus\{\overline{x}\}$. By Proposition˜3.15, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in U$ after possibly reducing $U$. Thus, for all $x\in U\setminus\{\overline{x}\}$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{x},r)<\accentset{\circ}{f}(x,r)$ for all $r\in(0,\overline{r}]$. ∎
Corollary 3.19.
@@
One argues as for Corollary˜3.19 using Proposition˜3.17. ∎
@@
Fact˜3.22 implies that $x\preceq y$ and $y\preceq x$ for all $x,y\in G\overline{x}$. By Proposition˜3.15, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in\mathbb{R}^{n}$. There exists a neighborhood $U$ of $G\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(x)<\mathrm{lip}f(y)$ for all $x\in G\overline{x}$ and $y\in U\setminus G\overline{x}$. Hence, for any such $x$ and $y$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(x,r)<\accentset{\circ}{f}(y,r)$ for all $r\in(0,\overline{r}]$. ∎
@@
Fact 3.25.
[44, Theorem 10.6] Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is locally Lipschitz regular and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. Then $f$ is regular and
@@
Remark 3.27.
Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is $C^{1}$ such that $\nabla g(0)=0$ and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is locally Lipschitz. The Jacobian chain rule [12, Theorem 2.6.6] yields $\overline{\partial}f(x)=\overline{\partial}F(x)^{T}\nabla g(F(x))$. Hence, if $F(x)=0$, then $\overline{\partial}f(x)=\{0\}$. Since the Clarke subdifferential is a singleton, by [44, Theorem 9.18] $f$ is strictly differentiable [44, Definition 9.17] at $x$, $\nabla f(x)=0$, and $\mathrm{lip}f(x)=0$.
@@
Fact 3.28.
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be lower semicontinuous and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n)$. For all $\overline{x}\in\mathrm{dom}f$ and $g\in G$, $\partial f(g^{-1}\overline{x})=g^{T}\partial f(\overline{x})$.
@@
Applying the chain rule [44, Exercise 10.7] to $f(x)=f(gx)$ at $g^{-1}\overline{x}$ for any $g\in G\subseteq\mathrm{GL}(n,\mathbb{R})$ yields the result. ∎
Fact 3.29.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n)$. For all $\overline{x},v_{1},\ldots,v_{k}\in\mathbb{R}^{n}$, $f^{(k)}(g^{-1}\overline{x})(v_{1},\ldots,v_{k})=f^{(k)}(\overline{x})(gv_{1},\ldots,gv_{k})$.